# An agent-based model of behaviour change calibrated to reversal learning data

### v1

@@ V Conclusion
In summary, we developed a simple yet extensible agent-based model of behaviour change that is mechanistically adequate to simulate reversal learning. We calibrated the parameters of our model from empirical behavioural data using approximate Bayesian computation and validated that the model reproduces behaviours observed in single and serial reversal learning experiments given those parameters. This work is a step towards developing computational models of behaviour change whose parameters are fully identifiable from actual behavioural data and are grounded in psychological theories. Such models are essential for understanding and predicting how behavioural changes affect emergent real-world phenomena with better accuracy, reliability, and practicality.

### v2

@@ V Conclusion
In summary, we developed a simple yet extensible microsimulation model of behaviour change that is mechanistically adequate to simulate reversal learning. We calibrated the parameters of our model from empirical behavioural data using approximate Bayesian computation and validated that the model reproduces behaviours observed in single and serial reversal learning experiments given those parameters. This work is a step towards developing computational models of behaviour change whose parameters are fully identifiable from actual behavioural data and are grounded in psychological theories. Such models are essential for understanding and predicting how behavioural changes affect emergent real-world phenomena with better accuracy, reliability, and practicality.
Code availability
The code used for this work is available at: https://github.com/rddelosreyes/abm-behaviour-change